OpenStudy (anonymous):

Find all real zeroes for the following function: g(x)=x^2+x+1

geerky42 (geerky42):

Zeroes is the values of x where \(g(x) = 0\) Try to use quadratic formula: \(x = \dfrac{-b\pm\sqrt{b^2-4ac}}{2a}\), from \(ax^2+bx+c=0\)